A sinking fund amounting to

​$16,000

is to be created by making payments at the beginning of every

6 months

for

5

years. Interest earned by the fund is

9.6%

compounded

semi-annually.

Determine the size of the periodic payment and compute the interest earned during the fourth payment interval. Verify your answer by checking the sinking fund schedule.

The size of the payment is

​$nothing.

​(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as​ needed.)

Complete the​ table, rounding each entry to the nearest cent.

​(Do not include the​ $ symbol in your​ answers.)

Payment Interval Number

Periodic Payment​ (Beginning)

Payment Interval Interest

i=4.8%

Increase in Fund

Balance at End of Payment Interval

0

​-

​-

​-

0

1

nothing

nothing

nothing

nothing

2

nothing

nothing

nothing

nothing

3

nothing

nothing

nothing

nothing

4

nothing

nothing
Answer #1

Schh PV = I Govo $ Semj annually 9.6.7. ta syn PV = p (4+0).* p= 16000 X 48 you - (1 + 4.81-lo loo P= 2051-996 11 2052 2052 €w 2052 Payment arel interest paid to porincipal remaing = 13370.368 x 4-8 = 641.778 low 1410.222 = 11960.146 Payment 4th = 20